# Extending the European Competence Framework for Quantum Technologies: new proficiency triangle and qualification profiles

### v1

@@ Abstract
With the increasing industrial relevance of quantum technologies (QTs), a new quantum workforce with special qualification will be needed. Building this workforce requires educational efforts, from short-term training to degree programs. In order to plan, map and compare such efforts, personal qualifications or job requirements, standardization is necessary. The European Competence Framework for Quantum Technologies (CFQT) provides a common language for QT education. The 2024 update to version 2.5 extends it with the new proficiency triangle and qualification profiles: The proficiency triangle proposes six proficiency levels for three proficiency areas, specifying knowledge and skills for each level. Nine qualification profiles show prototypical qualifications or job roles relevant to the quantum industry with the required proficiency, examples, and suggestions. This is an important step towards the standardization of QT education. The CFQT update is based on the results of an analysis of 34 interviews on industry needs. The initial findings from the interviews were complemented by iterative refinement and expert consultation, as well as by ML-driven analysis and mapping of job posts, revealing recommendations for formulating job posts.
@@ 1 Introduction
Modern Quantum Technologies (QTs), such as quantum computers, quantum sensors or quantum communication devices, are gaining increasing industrial relevance. Building the future quantum workforce requires not only intensive educational efforts such as training and study programs. QT education also needs a common language as well as standardization to make efforts and qualifications comparable: The European Competence Framework for Quantum Technologies (CFQT) [1] is the “reference framework for planning, mapping and comparing QT-related educational activities, personal qualification and job requirements”. This paper documents the update from version 2.0 to 2.5 and the advances in the new version.
@@
The CFQT aims to structure QT competences as well as related competences relevant to the future quantum workforce. It was already used, for example, to map course content [2], to categorize research [3], or as the backbone for a curriculum transformation framework [4]. Version 2.5 (2024) of the CFQT was published by the European Commission, EU Publications Office (see Ref. [1]). It was developed within the European Quantum Flagship [5] Coordination and Support Action (CSA) projects QTEdu [6] and QUCATS [7]. Within QUCATS, the European Quantum Readiness Center (EQRC) [8] has been launched to establish best practices, including standardization, using the CFQT as a tool.
@@
Version 2.0 of the CFQT focused on content relevant in the context of QT, structured in the content map. The content map and the details for the eight content domains have remained consistent from version 2.0 (2023) to 2.5 (2024). With the update, two key additions have been made: the proficiency triangle (see Sec. 3.1) and the qualification profiles (see Sec. 3.2). Appendix A.1 provides an overview of the general structure of the CFQT document.
@@
Crucial input for the CFQT extension in version 2.5 came from the analysis of 34 interviews focusing on QT educational needs in industry. The primary objective of the interview analysis was to report on the qualification and training needs in industry, as documented in Ref. [9]. However, the insights from the interviews provided great input for the CFQT update. In addition to the manual analysis, a GPT was employed to analyze anonymized interview transcripts and extract typical job roles to improve the initial qualification profiles. The GPT was also used to map job posts to the qualification profile descriptions and thus evaluate the improvement of the profile descriptions through three draft versions.
@@
The development of the CFQT started with an iterative study in 2020/2021 [10], leading to version 1.0 [11] compiled within the QTEdu CSA [6]. Within the QUCATS project, it was updated to version 2.0, as documented in Ref. [12]. The update to version 2.5, which brought the proficiency triangle and qualification profiles into the framework, was released in April 2024. The process and results of the update are documented below, including several excerpts from the CFQT document for illustration. A version history and related publications are given in Appendix A.2.
2.1 Interview analysis and iterative refinement
In summer 2023, 34 interviews with industry representatives were conducted and analyzed to answer research questions including What QT qualification and training needs are reported by industry? An in-depth analysis is documented in Ref. [9], including details about the methodology, which was based on qualitative content analysis [13]. The interviewees represent QT companies or companies with QT departments, from start-ups to very large companies, covering technical, management or business roles.
In some of the interviews, the material ‘competence types’ (Fig. 5) was shown to stimulate the flow of the interviews. This was an outline for structuring typical competences that we expected to be relevant in industry. Three levels were listed for each of the three ‘competence types’: build/develop QT (level B–Build…, C–Component improvement…, D–Develop…), overview/communicate about QT (level N, O, P), and use/adapt QT (U, V, W), see Appendix B. The key competences and qualifications discussed in the interviews were categorized according to this structure. However, it turned out that it was incomplete and not disjunctive, so refinement was needed.
@@ 2.1 Interview analysis and iterative refinement
Through an iterative process, this initial structure was refined into what we named the proficiency triangle, see Sec. 3.1. It consists of three proficiency areas with six proficiency levels each. For each of these 18 levels, a competence statement was formulated using action verbs [14]. They are complemented by descriptions of knowledge and skills whose formulations are based on the European Qualifications Framework (EQF) [15]. Already for the CFQT version 2.0 [16], the EQF was used to describe six proficiency levels, as documented in Ref. [12]. The proficiency level labels A1, A2, B1, B2, C1 and C2 are associated with the levels commonly used to specify qualifications in languages in Europe [17]. The DigCompEdu [18] framework was used as a template which also provided the initial version of the level keywords, e.g. awareness for level A1. Those were also updated to better match the level descriptions.
All proficiency levels were formulated in a generic way, adapted from the EQF definitions for QT in general. In the new version 2.5, they are specified for the three proficiency areas, and key competences identified in the interviews were incorporated.
In addition, categories related to job roles were formulated. During the qualitative content analysis of the interview transcripts, eleven categories were used, mainly based on the structure of the ‘competence types’ (Fig. 5). In Table 5, these job role categories are shown, together with a reference to the related ‘competence type’ level.
In the next step, 30 anonymized interview transcripts were analyzed using a large language model (LLM, namely GPT3.5) to identify frequently mentioned personas representing categories or clusters of job roles. This resulted in the descriptions of the 10 personas provided in Appendix C.1. These personas were checked against the initial job role categories, supporting the reduction from eleven to nine profile-related categories and an initial refinement. For the remaining categories, a second iteration of categorizing the interview passages was performed. Several categories in the second iteration are already named similarly to the final profiles, as documented in Table 5. Based on these categories, or more precisely, the categorized interview passages, the qualification profile descriptions version 2.1 were created.
These profiles were then refined through discussions with several colleagues and by gathering feedback from experts in both academia and industry. In this way, the titles and descriptions were refined, and example personas and suggestions were added based on the results of the interview analysis documented in Ref [9].
2.2 LLM-based evaluation of the qualification profiles
As a first trial of the consistency of the qualification profiles and the robustness of their descriptions, we used an LLM in order to test the profiles for stability under repeated classifications, adapting a methodology previously used for sentiment classification [19]. We used a dataset of 50 QT related job descriptions taken from a larger collection of LinkedIn job posts collected for research at the EQRC [8]. The job descriptions were classified into one of the nine qualification profiles using the LLM.
We ran the matching algorithm 15 times with the LLM gpt3.5-turbo-16k and calculated the stability of the classification as $\frac{x}{15}$, where $x$ is the most frequent classification into one profile for each job post, which is divided by the number of total classifications (15). For all 50 jobs, the mean value $s=\frac{1}{50}\sum\left(\frac{x}{15}\right)$ was calculated as a metric of the classification stability. For the three versions of the profile descriptions, these are $s_{1}=0.944$ (v2.1), $s_{2}=0.927$ (v2.2) and $s_{3}=0.981$ (v2.3). This demonstrates an increase from the first version to the third version. The full distribution of profile classifications is visualized in Fig. 6, data are available on Zenodo [20].
In all versions, the majority of the jobs were stable under 15 classifications, as measured by the stability $s$. However, some jobs were split across classes. For version 2.3, three jobs descriptions were less than 85% stable under repeated classifications (i.e., the most frequent classification $x$ was less than 13 within 15 repetitions, thus IDs 10, 37 and 40 in Fig. 6). In order to understand this, these job descriptions were examined individually. This has implications for how job posts should be written in order to most clearly communicate the role, which we consider in Sec. 3.3.
2.3 Limitations
The update of the CFQT is primarily based on the analysis of 34 interviews with participants from the quantum industry. This is a strength of the dataset, as it is directly constructed from interviews with experts in the field. However, we should anticipate and acknowledge that there may be a bias due to the group of interviewees, as discussed in Ref. [9]. Most of the perspectives included come from people who have a very positive perception of QT, both in the interviews and in the iterative refinement. Therefore, they may identify personas more readily than those without direct involvement in the industry would.
The extraction of personas from the interview transcripts by an LLM improves the objectivity of the qualitative content analysis by having a kind of second coder analyze most of the interviews (the 30 transcripts that could be anonymized) and prepare a second set of categories.
3 Results: Proficiency triangle, qualification profiles and implications from the job post mapping
3.1 Proficiency triangle: 3 proficiency areas with 6 levels each
The new proficiency triangle adds a second dimension, proficiency, to the content map from the CFQT version 2.0. It consists of three proficiency areas (I) quantum concepts, (II) QT hardware and software (HW & SW) engineering, and (III) QT applications & strategies. The proficiency areas are described in the CFQT document, for example:
The proficiency of an individual may grow in these three areas independently, however, typically someone working in the QT industry would have proficiency in all three areas. For each of these three proficiency areas, six proficiency levels A1, A2, B1, B2, C1 and C2 are specified. Together they form the proficiency triangle shown in Fig. 1.
With the coverage of the triangle, a QT-specific set of knowledge and skills, or a qualification can be visualized. The qualification profiles show prototypical coverage of the proficiency triangle: Each profile is visualized by a partially colored proficiency triangle, indicating the level of proficiency with respect to each of the three proficiency areas, see Sec. 3.2.
Proficiency level Area (I) Area (II) Area (III)
A1 Awareness Reproduce basic q. concepts & terminology Reproduce basic functionalities of a QT facet Recognize potential of QT
A2 Literacy Describe fundamental q. concepts Perform basic tasks on a QT facet Identify value of QT
In addition, for each of the proficiency levels, relations to the content map (sub)domains (5.6, 6.7, …) and examples of how the levels could be reached are given. For example, the description of the lowest proficiency level A1 for the proficiency area (III) QT applications & strategies is:
A1 Awareness: Recognize potential of QT
  • K:
  • S:
with the related content (sub)domains & examples:
Multiple choice questions on possibilities/limitations. [quoted from 1, p. 17]
Examples of how to test or measure whether someone has reached this level are included (in italics). They are a first step towards a certification scheme. The proficiency level descriptions are formulated independently of, e.g., a concrete QT. To specify, e.g., a specific qualification of a person or the aim of a course or the requirements for a job, the proficiency triangle must be combined with a selection from the content map. For each level, there are references to related content (sub)domains. In the example above these are:
  • 5.6
    applications of quantum computing and simulation, including material science, engineering and design, optimization …   [p. 10]
  • 6.7
    applications of quantum sensors, including metrology at single quantum level, medicine and molecular biology, …  [p. 11]
  • 7.4, 7.6
    applications of quantum cryptography …; system networks (composite systems), quantum internet applications, … [p. 12]
  • 8.2
    business strategy, entrepreneurship and management (business level), including risk management and strategic decision making, … [p. 13]
  • 8.4
    responsibility and awareness, including ethics, public communication and outreach, education … [p. 13]
Alternatively, the focus could be on the value and impact for society and education, i.e. subdomain 8.4 instead of 8.2. With 8.4 one can map qualifications outside of industry, e.g. a school teacher who recognizes the potential of QT as a context for teaching quantum concepts and who has an overview of how QT can impact society as a whole. This would be relevant, e.g., to discuss QT and their expected impact already in high school and thus to get more school students interested in quantum (and STEM). Raising this interest was rated as very necessary (40 of 52 participants rated ‘high need’, 11 rated ‘low need’, one rated ‘no need’) in our follow-up survey to the interview analysis, see Ref. [9, Sec. 5.4].
3.2 Qualification profiles
The qualification of a person or by a training can be described by a coverage of the proficiency triangle, which specifies the proficiency level for each of the three proficiency areas. Since these are formulated independently of the concrete subject matter, they have to be combined with a selection from the content map. Nine qualification profiles included in the CFQT version 2.5 show the prototypical coverage of the proficiency triangle. They are listed in Table 2 with the required proficiency levels for each of the three proficiency areas.
Table 2: Qualification profiles and related proficiency levels in the CFQT version 2.5.
Proficiency level in area
  (I)  (II) (III)
No. Profile title q. backgr. QT eng. appl./strat.
P1 QT aware person A1 A1 A1
P2 QT informed decision maker A1
P3 QT literate person (QT literate business role, advocator, enthusiast) A2 A2 A2
P4 QT practitioner (working with QT, technician, QT user) A1 B1 A1
P5 QT business analyst A2 A2 B2
P9 QT core innovator C2 B1 A2
For each profile, a detail page provides the partially colored proficiency triangle with the descriptions of the highest proficiency level covered for each of the three proficiency areas. The profile is described generally, and example personas provide more concrete descriptions of what a job with that qualification might look like, for example for profile 4:
General description
The QT practitioner is someone who works around the development, assembly and operation of QT (technicians or ‘classical’ engineers with some QT specific additional qualifications), or uses QT with some customization:
Example personas
‘Classical’ engineer with QT add-on, QT lab technician (e.g. for operation and maintenance), QT assembly and test technician, …
An engineer working on QT development, could be an electronic or mechanical engineer or a software engineer/computer scientist, working on the control hardware/software for a qubit, needs mainly traditional engineering skills, but works together with the quantum people, so needs an idea of the special challenges in QT development (but does not need to understand the details, has a supervisor who ensures compliance with quantum requirements) and also has an idea of the applications etc. to know what they are working for. […] [1, p. 22]
3.3 Additional findings and implications from the job mapping
As discussed in Sec. 2.2, 50 job posts were classified by an LLM to three versions of the nine qualification profile descriptions. Most job posts were stable under repeated classification, visible in the statistical distribution shown in the Appendix, Fig. 6. Table 3 lists the titles of the qualification profiles in the three versions used for the job post mapping and the published version.
Table 3: Evolution of titles of the qualification profiles. Although the titles may not have changed between two iterations, the descriptions may have changed significantly.
No. v2.1 v2.2 v2.3 v2.5 (published in Ref [1])
1 QT aware person
2 QT aware decision maker QT informed decision maker
3 QT-literate communicator QT communicator QT literate person
4 QT (market) analyst QT business analyst renumbered to P5
5 QT user, engineer working with QT QT technician P4 QT practitioner
7 senior QT engineer/QT architect QT (HW or SW) specialist
8 QT strategist QT (product) strategist
9 QT core innovator
We notice that the jobs matching to the qualification profiles are not evenly distributed. For example, there are more job offers matching to profiles 5 (renumbered to profile 4 in the published version, see Table 3), 6, and 7 – the more technical and engineering roles – than those matching to the others. This is consistent with research currently being conducted within the QUCATS project at the EQRC [8, subpage analysis], which indicates that the majority of jobs in the QT field are currently technical, such as researchers or engineers, rather than non-technical, such as in business or management.
Taking a closer look on unstable job posts, we identified that these contained descriptions that did not represent the role as accurately as those with more stable classifications, or did not represent them at all, rather describing the company. For example, these unstable job descriptions included a) a software developer (ID 40): a very long (760 words) description of a company with no reference to the tasks of the role itself; b) an algorithm developer (ID 37), consisting of a description of the role of an engineering staff (profile 6) but also emphasizing tasks for a more senior position (profile 7); c) an open call for applicants of any role (ID 10, mapping to profiles v2.3).
As a practical result of the job post mapping, we conclude that not all job descriptions are completely suited to be mapped using the nine qualification profiles. We recommend that job descriptions be written with at least three specific sections: (a) Skills required to be qualified to perform the job, (b) tasks/responsibilities of the role, and (c) formal qualifications required to apply. Out of 1113 QT job posts aggregated by the EQRC [8] since January 2023, 895 (80.4%) are written with all three of these recommended sections. Therefore, roughly one in five job posts in the field is written with insufficient information to reliably profile the role. In order for graduates to be able to prepare for the job market, with full knowledge of the skills required for the position they are applying for, we recommend that employers consider including complete information in their job posts, enabling the role to be clearly understood.
4 Discussion and conclusion
Prior to implementing new courses and programs to develop the workforce pipeline in QT, it is critical to answer a key question: What specific knowledge and skills are needed for industrx growth? Different studies were conducted to analyze these needs [21, 22, 23, 24, 25, 10], including our analysis of 34 interviews with industry representatives [9] This research provided the primary input for the extension of the CFQT with the proficiency triangle and qualification profiles. The proficiency levels organized in the proficiency triangle cover key knowledge and skills discussed in the interviews, and the qualification profiles are supplemented by example personas and needs and suggestions extracted from the interview analysis.
  • educators to plan educational activities and map their objectives;
  • learners to identify what training or course they should take;
Early users are educators [e.g. 2, 4]. Two EU-funded projects agreed to use the framework to map their activities: DigiQ [26], developing 16 QT master’s programs across Europe, and QTIndu [27], developing industrial training courses on QT.
I think it is always better to have a standard. So the question is how to standardize these trainings? […] It comes back to the ECTS systems from education. I mean, they are standards.  [quote from an interview with a QT start-up; 9, p. 24]
For the update to version 2.0, we already aligned the proficiency level descriptions with the EQF and thus the European Credit Transfer and Accumulation System (ECTS), as described in Ref. [12, p. 3]. Following the EQF [15, p. 16], the three highest EQF levels are associated with the first, second and third cycles of the Qualifications Framework of the European Higher Education Area, which are based on the ECTS. Therefore, as the CFQT levels are formulated based on the EQF. The level B2 is linked to bachelor, C1 to master and C2 to doctoral programs. However, it should be emphasized that this mapping is not necessarily one-to-one, and that a degree program is not the only way to reach these levels. Especially in the business-related area (III), the high levels are more likely to be achieved through years of experience.
In version 2.0 of the CFQT, these links were explicitly stated in the proficiency level descriptions. For example, it was noted for level B2: “e.g. through a short research project as for a bachelor thesis, internship with project” [16, p. 5]. As mentioned above, this brings the risk of assuming a study program as the only way to reach a level, while there is a variety of opportunities to gain such a qualification.
In version 2.5, such comments are less prominent and vary across the three proficiency areas. For the same proficiency level B2 Investigation, these are for area (I) “Short research project with documentation, e.g. student research project or bachelor thesis”, for area (II) “Write report on requirements analysis and performance results” and for area (III) “Document an analysis for a concrete potential use case covering potential advances, risks, […]” [1, p. 15-17]. With the update, the focus in these remarks shifts from an approximate duration to what an individual should be able to answer, perform or prepare for a qualification covering the corresponding proficiency level. Suggestions for training formats, the need for study programs or work experience are provided for the qualification profiles, rather than being fixed for a proficiency level.
The qualification profiles, together with the suggestions listed on the detail pages, are the starting point for a certification scheme to ensure comparability of industry training across Europe. This scheme is currently being developed within the Quantum Flagship coordination project, QUCATS [7]. As a part of QUCATS, the EQRC also highlights “accords” (best practices) of organizations across the EU [8, subpage Accords]. Within these accords, the CFQT is a tool that can provide structure and standardization. As a recent example, the accord submitted by the QuantUM Group at the University of Minho provides a profile of a software engineer in quantum computing as a representative graduate from the new master degree established there [8, subpage Accords: Educational Accords]. In addition, the EQRC website hosts a curated playlist of videos, sorted based on the content map of the CFQT [8, subpage Resources]. The CFQT, consisting of the content map and the proficiency triangle with associated qualification profiles, provides a reference framework for greater standardization across Europe. However, adoption is critical to its success. Industry representatives seeking new employees should consider using the CFQT to understand their needs. They should apply the CFQT to define skills, requirements and responsibilities in their job advertisements. Educational institutions also have an important role to play. Here, the qualification profiles can be a step towards the future of standardized learning outcomes in QTs.
Abbreviations
  • CFQT
    European Competence Framework for Quantum Technologies [1]
  • CSA
    Coordination and Support Action
  • DigiQ
    Digitally enhanced quantum technology master (project title) [26]
  • EQF
    European Qualifications Framework [15]
  • EQRC
    European Quantum Readiness Center [8]
  • GPT
    Generative Pre-trained Transformer (one type of LLMs)
  • HW
    Hardware
  • LLM
    Large Language Model
  • ML
    Machine Learning
  • QT/QTs
    Quantum Technology/-ies
  • QTEdu
    Quantum Technology Education (QTEdu CSA, project title) [6]
  • QTIndu
    Quantum Technologies courses for Industry (project title) [27]
  • QUCATS
  • STEM
    Science, Technology, Engineering, and Mathematics
  • SW
    Software
Declarations
[…]
Data availability The three versions v2.1-v2.3 of profile descriptions, the 50 job posts, the prompt and the mapping statistics are available on Zenodo, see Ref. [20]. Here, also the data from the industry needs analysis (Ref. [9]) is available.
[…]
Acknowledgements We thank Carolyn Levy, Dennis Green, Dion Timmermann, and all experts from industry and academia for their valuable feedback within the iterative refinement of the descriptions of the proficiency level or qualification profiles. FG prepared, conducted and analyzed the interviews, prepared the CFQT drafts, collected feedback and wrote the original manuscript. SG conducted the machine learning based analysis and wrote the related part of the original manuscript (Sec. 2.2 and 3.3). All authors contributed to the study design and analysis, i.e. they discussed the methods and results, revised the manuscript, and read and approved the final manuscript.
[…]
Appendix A Structure and version history
A.1 Principal structure of the CFQT
The CFQT [1] consists of three main parts:
  1. 1.
  2. 2.
    The proficiency triangle visualizes six proficiency levels for three proficiency areas (p. 14; Fig. 1). In-depth level descriptions and more are provided for each of the three proficiency areas (p. 15-17). See Sec. 3.1.
  3. 3.
    The qualification profiles show typical qualification, i.e. coverage of the proficiency triangle. An overview (p. 18) and detailed descriptions including example personas and (training) suggestions (p. 19–27) are available. See Sec. 3.2.
A.2 Version history and related publications
Date [Ref.] Title/version Comment
June 2022
(Nov. 2020)
Sep. 2021 [*] CFQT version 1.0 Only graphical update.
June 2023
(Aug. 2022)
July 2023 [12] Towards a quantum ready workforce: the updated CFQT Paper on the update to v2.0.
In some of the interviews (22 out of 34), the material ‘Competence types’ was shown, reproduced in Fig. 5. They formed the initial system that was refined and extended through the CFQT update process, as described in Sec. 2.1.
Figure 5: Competence types as a material shown in the interviews, June 2023, outdated. For better readability, the descriptions are reproduced below.
The three competence types shown in Fig. 5 with three levels each are:

### v3

@@ Abstract
With the increasing industrial relevance of quantum technologies (QTs), a new quantum workforce with special qualification will be needed. Building this workforce requires educational efforts, ranging from short term training to degree programs. In order to plan, map and compare such efforts, personal qualifications or job requirements, standardization is necessary. The European Competence Framework for Quantum Technologies (CFQT) provides a common language for QT education. The 2024 update to version 2.5 extends it by the new proficiency triangle and qualification profiles: The proficiency triangle proposes six proficiency levels for three proficiency areas, specifying knowledge and skills for each level. Nine qualification profiles show prototypical qualifications or job roles relevant to the quantum industry, with the required proficiency, examples, and suggestions. This is an important step towards the standardization of QT education. The CFQT update is based on the results of an analysis of 34 interviews on industry needs. The initial findings from the interviews were complemented by iterative refinement and expert consultation.
@@ 1 Introduction
Modern Quantum Technologies (QTs), such as quantum computers, quantum sensors or quantum communication devices, are gaining increasing industrial relevance. Building the future quantum workforce requires not only intensive educational efforts such as training and study programs. QT education also needs a common language as well as standardization efforts to make education and qualifications comparable: The European Competence Framework for Quantum Technologies (CFQT) [1] is the “reference framework for planning, mapping and comparing QT-related educational activities, personal qualification and job requirements”. This paper documents the update from version 2.0 to version 2.5 and the advances in the new version.
@@
The CFQT aims to structure QT competences as well as related competences relevant to the future quantum workforce. It was already used, for example, to map course content [2], to plan new courses [3], to categorize research [4], or as the backbone for a curriculum transformation framework [5]. Version 2.5 (2024) of the CFQT was published by the European Commission, EU Publications Office (see Ref. [1]). It was developed within the European Quantum Flagship [6] Coordination and Support Action (CSA) projects QTEdu [7] and QUCATS [8]. Within QUCATS, the European Quantum Readiness Center (EQRC) [9] has been launched to establish best practices, including standardization, using the CFQT as a tool.
The objective of the update to version 2.5 was to provide more detailed and substantiated descriptions for the proficiency levels, as well as a new version for the qualification profiles from 2022 [beta version, 10]. General descriptions for six proficiency levels were already added in the CFQT update to version 2.0, but needed further refinement and specification for QT – as in the new proficiency triangle (see Sec. 3). The proficiency triangle enables the specification of proficiency or qualification in three dimensions, covering (I) quantum concepts, thus physics fundamentals, (II) QT engineering competences, including QT functionalities, and (III) QT applications and strategies, i.e. the business perspective, including aspects such as impact and ethics. By assigning a proficiency level for each of these three proficiency areas, the qualification of an individual can be specified, for example.
@@
A total of nine new qualification profiles were identified and visualized with the proficiency triangle (see Sec. 4). They represent prototypical qualifications relevant to industry. However, to really specify such a qualification, a second measure is needed: A selection from the content map. It structures content relevant in the context of QT and was in the focus of the update to version 2.0 [11]. The content map and the details for the eight content domains remained consistent from version 2.0 (2023) to 2.5 (2024). In the new version, two examples are given of how to combine such a selection from the content map and a profile based on the proficiency triangle. Appendix A.1 provides an overview of the general structure of the CFQT document, from the content map and the proficiency triangle to the qualification profiles and examples.
@@
Crucial input for the CFQT extension in version 2.5 came from the analysis of 34 interviews focusing on QT educational needs in industry. The primary objective of the interview analysis was to report on the qualification and training needs in industry, as documented in Ref. [12]. However, the insights from the interviews provided great input for the CFQT update: key competences were identified and incorporated into the proficiency level descriptions, and the jobs and roles discussed in the interviews led to the first draft of the new qualification profiles. In addition to the manual analysis, a GPT was employed to analyze anonymized interview transcripts and extract typical job roles to improve the initial qualification profiles.
@@
The development of the CFQT started with an iterative study in 2020/2021 [13], leading to version 1.0 [14] compiled within the QTEdu CSA [7]. Within the QUCATS project [8], it was updated to version 2.0, as documented in Ref. [11]. In this update, the content map was revised and descriptions of six proficiency levels A1 to C2 were added. The proficiency level descriptions were based on the European Qualifications Framework (EQF) [15], thus specifying the required knowledge and skills for each proficiency level adapted to QT in general.
@@
The proficiency level labels A1, A2, B1, B2, C1 and C2 are associated with the levels commonly used to specify qualifications in languages in Europe [16]. The DigCompEdu [17] framework was used as a template which also provided the initial version of the level keywords, e.g. awareness for level A1. These keywords were also updated in version 2.5 to better match the level descriptions.
The update to version 2.5, which brought the proficiency triangle and qualification profiles into the framework, was released in April 2024. The process and results of the update are documented below, including some excerpts from the CFQT document for illustration. A version history and related publications are given in Appendix A.2.
2.1 About the interviews: introduction of ‘competence types’
In summer 2023, 34 interviews with industry representatives were conducted and analyzed to answer research questions including What QT qualification and training needs are reported by industry? An in-depth analysis is documented in Ref. [12], including details about the methodology, which was based on qualitative content analysis [18]. The interviewees represent QT companies or companies with QT departments from different European countries, from start-ups to very large companies, covering technical, management or business roles.
In some of the interviews, a slide about the ‘competence types’ (reproduced in Fig. 7) was shown to stimulate the flow of the interviews – others were conducted without this input, e.g. if the interviewee expressed very clear or different views in the previous questions. This material introduced a structure of competence descriptions for three levels for each of the three ‘types’ overview, communicate (level N, O, P), build, develop (level B, C, D) and use, adapt (level U, V, W):